OpenStudy (baldymcgee6):

Which formula based on partial derivatives provides the slope of the level curve z = f(x,y) ?

OpenStudy (baldymcgee6):

I know the answer... But I don't know why.

zepdrix (zepdrix):

What is the answer? Maybe it'll refresh my memory ^^ heh

OpenStudy (baldymcgee6):

I cant get the del sign in latex :/

zepdrix (zepdrix):

its ummm nambla i think.. \nambla

OpenStudy (baldymcgee6):

i thought it was \del

zepdrix (zepdrix):

nabla*

OpenStudy (baldymcgee6):

i'll try :)

OpenStudy (baldymcgee6):

nope thats an upside down triangle

zepdrix (zepdrix):

Yah that's the... del operator :o do you want the one with the vector arrow above it..? :o

OpenStudy (baldymcgee6):

\frac{ dy }{ dx } = -\frac{\del f /\del x}{\del f/\del y}

OpenStudy (baldymcgee6):

i mean like the one used on http://en.wikipedia.org/wiki/Partial_derivative

OpenStudy (baldymcgee6):

\[\partial \]

OpenStudy (baldymcgee6):

ahhh \partial

OpenStudy (baldymcgee6):

\[\Huge\frac{ dy }{ dx } = -\frac{\partial f /\partial x}{\partial f/\partial y}\]

zepdrix (zepdrix):

\[\huge \vec \nabla f(x,y)=\left<\frac{\partial f}{\partial x}, \frac{\partial f}{\partial y}\right>\]

zepdrix (zepdrix):

Oh sorry i didn't notice u already found it, hehe

OpenStudy (baldymcgee6):

oo, yours is fancy, i dont know what the nabla is though, thats okay...

zepdrix (zepdrix):

Yah i been trying to figure out latex the last couple of weeks :) it's kinda interesting

OpenStudy (baldymcgee6):

very interesting.. i'm pressured to use it for one of my math classes, i try to avoid it.. :)

OpenStudy (baldymcgee6):

anyways, do you get the question?

zepdrix (zepdrix):

the nabla thing is the "gradient" of f, it's the direction of greatest increase from a surface. It ummmmmm... i would try to draw a picture but i'm not sure i have the greatest understanding of it myself... :\ hmm

OpenStudy (baldymcgee6):

haha, thats okay.. i'll learn that when i get to that.